OpenStudy (anonymous):

What is the solution set of this system of equations? y = x2 − 3x − 4 x = y + 8

OpenStudy (anonymous):

Flip the first second equation so you have the \(y\)s on the same side.\[ y+8=x \]

OpenStudy (anonymous):

Then you subtract the second one from the first. \[ \underbrace{y}_{-(y+8)} = \underbrace{x^2-3x-4}_{-x} \]

OpenStudy (anonymous):

You will get: \[ -8=x^2-4x-4 \]This is a quadratic equation. Can you solve it now?

OpenStudy (anonymous):

yes, thank you